# Disorder-Engineered Hybrid Plasmonic Cavities for Emission Control of Defects in hBN

### v1

@@ II.1 AgNPs on hBN: Emitter Characterization and Plasmonic Coupling
The saturation profile of the ZPL emission reveals three significant changes upon coupling to the small AgNPs. First, under weak excitation conditions, the ZPL intensity exhibits a quenching factor of approximately 8.4, indicating a strong reduction in the quantum efficiency of the emitter. This low-power behavior arises from a combination of field enhancement and modifications to the emitter’s decay dynamics. Specifically, the enhancement of the non-radiative decay rate outweighs enhancements in both the excitation and radiative decay rates, a well-established effect for small AgNPs near quantum emitters [48]. Second, the excitation power required to reach the saturation regime increases by nearly a factor of 6 af ter AgNP coupling. This is consistent with an increase in the total decay rate, as saturation intensity is reached when the excitation rate $\gamma_{12}$ becomes comparable to the sum of the decay channels. The AgNP-induced modification accelerates the decay processes, necessitating higher excitation power to achieve saturation. Finally, at excitation powers well above saturation, the ZPL intensity still shows a $\sim 40\%$ reduction compared to the uncoupled case. Since the saturation intensity is primarily determined by the radiative decay rate, this observation implies that coupling to the AgNP leads to a reduced radiative decay rate. Together, these results demonstrate that the presence of a small AgNP strongly modifies both radiative and non-radiative processes, ultimately diminishing the quantum efficiency and emission intensity of the defect.
@@
Complementary FDTD simulations further support these findings by calculating radiative and non-radiative decay rates as a function of emitter–nanoparticle separation and particle size. As shown in Figure 3d, these simulations illustrate how a nearby AgNP (with a diameter of 40 nm) alters the emitter’s decay dynamics. The shaded regions in the plots represent the range of rate modifications arising from the orientation of the emitter’s dipole moment relative to the nanoparticle surface, ranging from a dipole aligned parallel to the nanoparticle surface (strongest coupling) to one oriented orthogonally (weakest coupling). This variation establishes upper and lower bounds for both radiative and non-radiative decay rate enhancements. At emitter–nanoparticle separations less than 10 nm, strong non-radiative enhancement dominates due to energy transfer into lossy plasmon modes of the metal, leading to fluorescence quenching. These effects are more pronounced for dipoles oriented orthogonally to the nanoparticle surface, which couple less efficiently to radiative modes and more strongly to non-radiative channels. Overall, the simulation results are in good agreement with experimental observations of lifetime shortening and emission quenching.
Figure 4: (a) PL Spectra of a defect with ZPL at 616 nm before and after the metal nanoparticle taken at similar excitation conditions. Broad peaks around 670 nm are the corresponding PSBs of the defect emission. (b) Excitation power-dependent ZPL intensity before and after the fabrication of AgNPs. A strong fluorescence enhancement is observed. (c) The result of time-resolved measurements on the ZPL emission before and after the AgNPs shows a strong decay time enhancement. (d) Results of FDTD simulations on an ideal system where a defect in hBN is coupled to a small AgNP with a diameter of 120 nm. Modifications on both radiative and non-radiative rates as a function of emitter-AgNP distance are clearly visible.
@@
The FDTD simulations, shown in Figure 4d, indicate that at emitter–NP separations of approximately 10–25 nm, the radiative decay rate is significantly enhanced, while non-radiative rates remain comparatively low. This favorable regime supports efficient photon emission and aligns well with the experimental PL and lifetime results. Additionally, the enhancement exhibits strong dependence on dipole orientation: emitters with dipole moments parallel to the nanoparticle surface couple more efficiently to the plasmonic mode, resulting in greater Purcell enhancement compared to orthogonal orientations.
@@
These results establish a clear nanoparticle size-dependent behavior in the plasmonic coupling of hBN quantum emitters. Engineering the size and geometry of metal nanoparticles thus offers a powerful means to tailor light–matter interactions in two-dimensional quantum photonic systems. Moreover, the observed changes in decay dynamics suggest that plasmonic coupling not only modifies radiative and non-radiative decay pathways, but also perturbs metastable triplet states. This can alter excited-state population dynamics and saturation behavior, consistent with plasmon-mediated control of excited-state kinetics reported in other solid-state emitters [52].
@@
II.2 Hybrid Nanocavities on Gold/Silicon Dioxide: Enhanced Uniformity and Confinement
Figure 5: (a) Simulated structure (not to scale), images of hBN without plasmonic nanocavity on the (b) conventional bright-field microscope and (c) confocal laser scanning microscope (CLSM), images of hBN with plasmonic nanocavity on the (d) conventional bright-field microscope, and (e) CLSM. The spectrum of hBN defect emissions at different wavelengths (f-i) without and (j-m) with plasmonic nanocavity.
@@ II.2 Hybrid Nanocavities on Gold/Silicon Dioxide: Enhanced Uniformity and Confinement
In the second phase of this study, building upon the emission enhancement observed via thermally dewetted AgNPs, we developed a hybrid cavity platform to further improve the enhancement of emission from defects in hBN nanoflakes. This platform consists of a metal-dielectric Au/SiO2 bilayer substrate combined with self-assembled AgNPs. Two parameters of the hybrid cavity platform, SiO2 and Au layer thicknesses, are investigated by FDTD simulations, and the thickness values of 20 nm and 100 nm are determined for SiO2 and Au mirror layer for the fabrication process, respectively. Scattering cross-section simulation results for the hybrid cavity with different thickness values can be found in the Supporting Information, Figures S2 and S3.
@@
Figure 5 presents (a) simulated structure (not to scale), (b) and (d) conventional bright-field microscope images in panels, and (c) and (e) fluorescence imaging in panels for samples without and with AgNPs, respectively. Panels 5 (f-i) show the PL counts from hBN flakes prior to AgNP formation. After the thermal dewetting process and formation of AgNPs, the PL intensities increased by approximately 100-fold, as shown in panels (j-m). Four distinct measurements exhibiting emissions with consistent spectral features were compared in this study. These emissions arise from randomly selected emitters, since tracking the exact same emitter before and after dewetting was not feasible.
@@ III Conclusion
This study establishes a powerful and scalable platform for engineering light–matter interactions at the nanoscale by integrating stochastic plasmonic nanocavities with quantum emitters in hBN. Through thermal dewetting of noble metal films, we demonstrate a deterministic-free fabrication route that yields strong and tunable coupling by-design between defect centers and plasmonic nanoantennas achieving photoluminescence enhancements of up to two orders of magnitude. This result firmly establishes these hybrid structures as efficient single photon sources for applications in quantum technologies, i.e. quantum key distribution [53, 54, 55, 56]. Our results reveal how local field confinement, decay rate manipulation, and cavity-emitter synergy can be harnessed without complex nanofabrication or precise emitter positioning.
Beyond the demonstrated optical control, this platform paves the way for transformative advances in quantum photonic technologies. Its simplicity, material compatibility, and spectral versatility make it a promising candidate for integration on next-generation on-chip quantum circuits and room-temperature nanoscale quantum sensors as tunable and bright single-photon sources. In particular, the ability to construct high-efficiency quantum nanoantennas on 2D materials such as hBN opens up unexplored opportunities for label-free single-molecule detection, quantum-enhanced sensing, and hybrid photonic-bio interfaces [57].
IV Experimental Section
@@
Multilayer hBN flakes were obtained commercially in solution form (Graphene Supermarket) and drop-casted (${\approx 10\ \mu}$L) on a silicon substrate, which has laser engraved marks to locate specific emitters before and after the fabrication of AgNPs.
@@
Supporting Information
Supporting Information is available from the Wiley Online Library or from the author.
Acknowledgments
The authors thank Selcuk Yerci for the workstation on which we performed our computations and Vasilis Karanikolas for reviewing some of the calculations. S. G. acknowledges the Middle East Technical University Postdoctoral Research Program (DOSAP-B). S. A. acknowledges the support of the Türkiye Scientific and Technological Research Council (TÜBİTAK) under project number 118F119.

### v2

@@ 2 Results and discussion
The saturation profile of the ZPL emission reveals three significant changes upon coupling to the small AgNPs. First, under weak excitation conditions, the ZPL intensity exhibits a quenching factor of approximately 8.4, indicating a strong reduction in the quantum efficiency of the emitter. This low-power behavior arises from a combination of field enhancement and modifications to the emitter’s decay dynamics. Specifically, the enhancement of the non-radiative decay rate outweighs enhancements in both the excitation and radiative decay rates, a well-established effect for small AgNPs near quantum emitters Lakowicz 2005. Second, the excitation power required to reach the saturation regime increases by nearly a factor of 6 after AgNP coupling. This is consistent with an increase in the total decay rate, as saturation intensity is reached when the excitation rate $\gamma_{12}$ becomes comparable to the sum of the decay channels. The AgNP-induced modification accelerates the decay processes, necessitating higher excitation power to achieve saturation. Finally, at excitation powers well above saturation, the ZPL intensity still shows a $\sim 40\%$ reduction compared to the uncoupled case. Since the saturation intensity is primarily determined by the radiative decay rate, this observation implies that coupling to the AgNP leads to a reduced radiative decay rate. Together, these results demonstrate that the presence of a small AgNP strongly modifies both radiative and non-radiative processes, ultimately diminishing the quantum efficiency and emission intensity of the defect.
@@
Complementary FDTD simulations further support these findings by calculating radiative and non-radiative decay rates as a function of emitter–nanoparticle separation and particle size. As shown in Figure 3d, these simulations illustrate how a nearby AgNP (with a diameter of 40 nm) alters the emitter’s decay dynamics. The shaded regions in the plots represent the range of rate modifications arising from the orientation of the emitter’s dipole moment relative to the nanoparticle surface, ranging from a dipole aligned parallel to the nanoparticle surface (strongest coupling) to one oriented orthogonally (weakest coupling). This variation establishes upper and lower bounds for both radiative and non-radiative decay rate enhancements. At emitter–nanoparticle separations less than 10 nm, strong non-radiative enhancement dominates due to energy transfer into lossy plasmon modes of the metal, leading to fluorescence quenching. These effects are more pronounced for dipoles oriented orthogonally to the nanoparticle surface, which couple less efficiently to radiative modes and more strongly to non-radiative channels. Overall, the simulation results are in good agreement with experimental observations of lifetime shortening and emission quenching. Detailed representative simulation environment is given in Supporting Information Figure S2.
To further elucidate the origin of fluorescence quenching observed for smaller AgNPs, wavelength-dependent absorption and scattering cross sections were calculated using FDTD simulations. As shown in the Supporting Information (Fig. S4), absorption dominates over scattering for small nanoparticles across the spectral range relevant to the hBN ZPL emission, explaining the enhanced non-radiative decay channels and reduced quantum efficiency.
Figure 4: (a) PL Spectra of a defect with ZPL at 616 nm before and after the metal nanoparticle taken at similar excitation conditions. Broad peaks around 670 nm are the corresponding PSBs of the defect emission. (b) Excitation power-dependent ZPL intensity before and after the fabrication of AgNPs. A strong fluorescence enhancement is observed. (c) The result of time-resolved measurements on the ZPL emission before and after the AgNPs shows a strong decay time enhancement. (d) Results of FDTD simulations on an ideal system where a defect in hBN is coupled to a large AgNP with a diameter of 120 nm. Modifications on both radiative and non-radiative rates as a function of emitter-AgNP distance are clearly visible. Simulation parameters were chosen to represent experimentally relevant nanoparticle sizes and emitter–nanoparticle separations; full details are provided in the Supporting Information.
@@
The FDTD simulations, shown in Figure 4d, indicate that at emitter–NP separations of approximately 10–25 nm, the radiative decay rate is significantly enhanced, while non-radiative rates remain comparatively low. This favorable regime supports efficient photon emission and aligns well with the experimental PL and lifetime results. Additionally, the enhancement exhibits strong dependence on dipole orientation: emitters with dipole moments parallel to the nanoparticle surface couple more efficiently to the plasmonic mode, resulting in greater Purcell enhancement compared to orthogonal orientations. These results establish a clear nanoparticle size-dependent behavior in the plasmonic coupling of hBN quantum emitters. Engineering the size and geometry of metal nanoparticles thus offers a powerful means to tailor light–matter interactions in two-dimensional quantum photonic systems. Moreover, the observed changes in decay dynamics suggest that plasmonic coupling not only modifies radiative and non-radiative decay pathways, but also perturbs metastable triplet states. This can alter excited-state population dynamics and saturation behavior, consistent with plasmon-mediated control of excited-state kinetics reported in other solid-state emitters Wang et al. 2020. The FDTD simulations presented here, and given in detail in Supporting Information Figure S3, are intended to capture the dominant physical trends governing plasmon–emitter interactions under experimentally relevant conditions, rather than to provide a one-to-one quantitative match to a specific emitter. Simulation parameters were therefore chosen to reflect the experimentally measured nanoparticle sizes and realistic emitter–nanoparticle separations, while accounting for the unknown dipole orientation of individual defects.
For larger AgNPs, the situation is markedly different. The corresponding wavelength-dependent cross sections (Supporting Information, Fig. S5) reveal a strong scattering contribution relative to absorption, which favors radiative decay-rate enhancement. This scattering-dominated regime is consistent with the experimentally observed increase in emission intensity and supports nanoparticle size as a key design parameter for engineering radiative versus non-radiative decay pathways.
The contrasting decay-rate modification observed for small and large AgNPs can be understood by considering the spectral overlap between the emitter’s ZPL and the wavelength-dependent absorption and scattering cross sections of the nanoparticles. In the strong coupling regime, where the emitter–nanoparticle separation is small, the dominant decay pathway is governed by whether the ZPL overlaps primarily with absorption-dominated or scattering-dominated plasmonic modes. If the ZPL lies within the absorption band, energy transfer into lossy plasmon modes enhances non-radiative decay and leads to fluorescence quenching. In contrast, spectral overlap with a scattering-dominated response favors radiative decay-rate enhancement, resulting in increased emission intensity. This framework provides a unified interpretation of the quenching observed for smaller AgNPs (Figure 3) and the enhancement observed for larger nanoparticles (Figure 4).
The FDTD simulations shown in Fig. 4d were performed using geometrical and material parameters selected to reflect experimentally relevant conditions rather than to reproduce a specific emitter–nanoparticle configuration. In the simulations, AgNPs with diameters matching the experimentally observed size distributions were coupled to an electric point dipole embedded in an hBN layer, with emitter–nanoparticle separations swept over the near-field interaction regime. The influence of dipole orientation was accounted for by considering both parallel and orthogonal dipole alignments. Detailed simulation parameters, including geometry, material models, boundary conditions, and numerical settings, are provided in the Supporting Information.
@@
Hybrid Nanocavities on Gold/Silicon Dioxide: Enhanced Uniformity and Confinement
Figure 5: (a) Schematic (not to scale) of the hybrid cavity geometry, indicating silver nanoparticles, hBN flakes, SiO2 spacer layer (20 nm), gold mirror layer (100 nm), and Si substrate. (b, d) Conventional bright-field microscope images with emitter regions marked by red circles (before AgNP formation) and blue circles (after AgNP formation). (c, e) Corresponding confocal laser scanning microscope (CLSM) images; red/blue circles indicate locations of spectra shown in panels (f–m). The apparent distortion in panel (e) is due to slight scanning calibration drift. (f–i) Representative PL spectra from emitters before AgNP formation. (j–m) Representative PL spectra from emitters after AgNP formation.
@@
In the second phase of this study, building upon the emission enhancement observed via thermally dewetted AgNPs, we developed a hybrid cavity platform to further improve the enhancement of emission from defects in hBN nanoflakes. The choice of a SiO2 spacer layer in the hybrid cavity architecture was deliberate and based on optical and practical considerations rather than substrate availability. SiO2 provides optical transparency in the visible range, chemical stability, and a moderate refractive index that enables controlled tuning of the emitter–mirror separation and the local density of optical states. The spacer thickness was optimized through FDTD simulations to balance near-field enhancement and non-radiative losses, making the Au/SiO2 configuration a common and widely adopted platform in plasmonic cavity designs. This platform consists of a metal-dielectric Au/SiO2 bilayer substrate combined with self-assembled AgNPs. Two parameters of the hybrid cavity platform, SiO2 and Au layer thicknesses, are investigated by FDTD simulations, and the thickness values of 20 nm and 100 nm are determined for SiO2 and Au mirror layer for the fabrication process, respectively. Scattering cross-section simulation results for the hybrid cavity with different thickness values can be found in the Supporting Information, Figures S6 and S7. Although thicker SiO2 spacer layers can yield higher absolute scattering cross sections for specific nanoparticle diameters, the experimentally observed emission enhancement occurs predominantly in the 550–650 nm wavelength range. Analysis of the wavelength-dependent absorption and scattering cross sections shows that a SiO2 thickness of 20 nm provides consistent spectral overlap between the hBN zero-phonon line emission and scattering-dominated plasmonic responses across all nanoparticle sizes investigated (70–90 nm).
It is important to note that tuning the spacer thickness does not represent a trivial optimization of the cavity response to match a specific emission wavelength. The emission spectrum of hBN defect centers is not tunable and exhibits significant emitter-to-emitter variability. Selecting a spacer thickness to maximize scattering at a single wavelength would therefore favor only a limited subset of emitters while compromising broadband coupling and uniformity. The 20 nm spacer thickness was thus chosen to ensure robust enhancement across multiple nanoparticle sizes and emitters, rather than to maximize the response for a single geometry.
@@
For comparison, the AgNP-on-hBN/Si configuration discussed earlier (Fig. 3 and Fig. 4) provides one-to-one emitter measurements, illustrating quenching and enhancement through controlled nanoparticle–emitter interactions, whereas the hybrid cavity geometry incorporates an additional Au/SiO2 mirror–spacer architecture that enables stronger field confinement, improved photon extraction, and enhanced uniformity across multiple emitters (Fig. 5), which are essential for scalable device concepts. As shown by the representative spectra in Fig. 5(f–m), different emitters exhibit ZPLs at distinct wavelengths, highlighting the intrinsically multispectral nature of defect emission in hBN and the broadband response of the hybrid cavity platform.
The reported $\sim$100-fold enhancement should therefore be interpreted as an indicative measure of the broadband enhancement capability of the hybrid cavity platform. It reflects the combined effects of plasmonic near-field coupling, optical feedback from the reflective Au layer, and the stochastic formation of AgNPs, rather than optimized coupling to a single, deterministically positioned defect.
@@ Conclusion
This study establishes a powerful and scalable platform for engineering light–matter interactions at the nanoscale by integrating stochastic plasmonic nanocavities with quantum emitters in hBN. Through thermal dewetting of noble metal films, we demonstrate a deterministic-free fabrication route that yields strong and tunable coupling by-design between defect centers and plasmonic nanoantennas achieving photoluminescence enhancements of up to two orders of magnitude. This result firmly establishes these hybrid structures as efficient single photon sources for applications in quantum technologies, i.e. quantum key distribution Wang et al. , Samaner et al. 2022, Al-Juboori et al. 2023, Tapşın et al. . Our results reveal how local field confinement, decay rate manipulation, and cavity–emitter synergy can be harnessed without complex nanofabrication or precise emitter positioning.
Beyond the quantum photonics implications demonstrated here, the broadband enhancement observed in our hybrid nanocavity platform could be leveraged in other domains. Plasmonic near-field coupling in the presence of nanoscale analytes has been shown to enable label-free single-molecule detection via enhanced light–matter interaction Herkert et al. 2023. The high field confinement and broadband spectral response also align with requirements for quantum-enhanced sensing schemes Dana et al. 2023. Furthermore, the solution-processable nature of our approach facilitates integration with biomaterials and microfluidic architectures, creating a pathway towards hybrid photonic–bio interfaces Wang et al. 2023. Such properties could also be harnessed for emerging paradigms such as single-photon-controlled neural networks Mousavi et al. 2023 and optogenetic stimulation Bansal et al. 2023, where precise, low-energy optical control is essential. While these applications were not experimentally demonstrated in this work, the observed optical enhancements and fabrication scalability suggest their feasibility.
While the present work focuses on a proof-of-principle demonstration, we note that the long-term stability of thermally dewetted AgNPs can be affected by oxidation. Literature reports indicate that dielectric overcoating or encapsulation methods can mitigate these effects and preserve plasmonic performance over extended periods Rycenga et al. 2011, Betancourt et al. 2021. Such strategies may be considered in future studies aimed at practical device implementations.
3 Methods
@@
Multilayer hBN flakes were obtained commercially in solution form (Graphene Supermarket) and drop-casted ($\approx 10\ \mu$}L) on a silicon substrate, which has laser engraved marks to locate specific emitters before and after the fabrication of AgNPs. The hBN flakes used in this study were commercially obtained and therefore contain naturally occurring intrinsic defect-based emitters. We did not employ defect creation steps such as ion implantation, irradiation, or high-temperature activation anneals. In this work, the term “defect” refers to optically active color centers naturally present in the as-received hBN flakes, which are subsequently coupled to plasmonic nanostructures.
To investigate the effect of Ag nanoantennas on the optical properties of single defects in hBN, the antennas were fabricated in the vicinity of defect sites using a self-organized solid-state dewetting approach, rather than relocating the emitters to pre-patterned structures Nguyen et al. 2018. For this purpose, Ag thin films of 5 or 13 nm were deposited on hBN-coated silicon substrates and subsequently annealed at $\sim 350,^{\circ}$C in a nitrogen environment Jiran and Thompson 1990, Borra et al. 2014, Nasser et al. 2013. This method is gentle on both the hBN flakes and the silicon substrate, which are stable up to $\sim 800\,^{\circ}$Kianinia et al. 2017. Silver was selected owing to its low optical losses, narrow plasmonic resonances, and favorable dewetting characteristics compared to other noble metals such as gold. The initial film thickness was used to tune the antenna size distribution, yielding average diameters of $\sim$35 nm and $\sim$110 nm for 5 nm and 13 nm films, respectively (Figure 2b,c). These regimes were chosen to probe both fluorescence quenching and enhancement in single hBN defects, since Ag nanoantenna dimensions strongly determine their absorption and scattering properties Lakowicz 2005. By controlling the dewetting duration, the size distribution of the resulting nanoantennas can be tuned (Supporting Information Figure S1). During annealing, the Ag film destabilizes well below the bulk melting point due to its high surface-to-volume ratio Jiran and Thompson 1992. Diffusion and agglomeration of Ag atoms lead to the spontaneous formation of hemispheroidal nanoislands distributed across both hBN flakes and the silicon substrate (Figure 2a). Owing to the stochastic nature of dewetting, a fraction of these nanoantennas form in close proximity to defect centers, enabling emitter–antenna coupling without additional positioning steps.
@@
For each substrate configuration, background spectra were recorded from adjacent hBN-free regions on the same substrates (including Au/SiO2 $\pm$ AgNPs) under identical acquisition settings and subtracted from emitter spectra prior to calculating ZPL-based enhancement factors; Raman and SERS features were excluded from the ZPL-based analysis. In our study, the “nanoparticle-only” condition is represented by AgNPs on hBN/Si (i.e., without the Au/SiO2 mirror–spacer), which isolates the near-field plasmonic effect and results in either modest enhancement (for large AgNPs) or fluorescence quenching (for small AgNPs), in agreement with expectations. A separate “hBN on Au/SiO2 without AgNPs” dataset is not included, as under our excitation and collection conditions stable single-emitter PL on Au/SiO2 prior to dewetting was not reliably detectable, indicating that the mirror–spacer alone does not account for the strong enhancement reported here. Likewise, “Au/SiO2 + AgNPs without hBN” cannot produce defect-related ZPL or PSB features; instead it serves to gauge background/SERS.
Author Contributions
S.G. and O.Y. contributed equally. S.G., O.Y., and A.B. conceived the study. S.G. and O.Y. performed the experiments and analyzed the data. C.R.-F., A.Y., H.C., and S.A. contributed to sample preparation, measurements, and interpretation. S.G., F.A., and A.Y. contributed to simulations. A.B., H.C., and S.A. supervised the project together. S.G., A.B., H.C., and S.A. wrote and reviewed the manuscript and all authors approved the final version.
Funding Sources
Türkiye Scientific and Technological Research Council (TÜBİTAK), project 118F119. Middle East Technical University Postdoctoral Research Program (DOSAP-B). Academy of Finland Flagship Programme (PREIN) (320165). Horizon2020 Marie SkłodowskaCurie grant agreement No. 895369.
{acknowledgement}
The authors thank Selçuk Yerci for the workstation on which we performed our computations, Vasilis Karanikolas for reviewing some of the calculations, and Melike Şeyma Tuna for help with the figures. S. G. acknowledges the Middle East Technical University Postdoctoral Research Program (DOSAP-B). S. A. acknowledges the support of the Türkiye Scientific and Technological Research Council (TÜBİTAK) under project number 118F119. H.C. acknowledges the financial support of the Academy of Finland Flagship Programme (PREIN) (320165). C.R.-F. acknowledges the Horizon 2020 research and innovation program under the Marie SkłodowskaCurie grant agreement No. 895369.
{suppinfo}
Annealing process, FDTD simulation details, Wavelength-dependent absorption and scattering cross sections, Spectral overlap between hBN ZPL and nanoparticle cross sections.